# Exploring Community-Powered Conversational Agent for Health Knowledge Acquisition: A Case Study in Colorectal Cancer

### v1

@@ 8. Discussion
In this paper, we have designed, developed, and evaluated CanAnswer that leverages community data for supporting users in learning about colorectal designs. Our formative study reveals that online communities are important sources for people to gain informational and emotional support about certain diseases, inline with previous work that highlights the benefits of online communities (Ding et al., 2020; Burgess et al., 2023; Sharma and De Choudhury, 2018; Sinha et al., 2018; Kendal et al., 2017). However, the unstructured and/or unauthoritative nature of the content in these communities could be troublesome (see Section 3.4.2). We developed a computational workflow (Figure 2) to identify and structure the useful content from a professional online community Dingxiang Doctor and a peer-led community RED, which enables our conversational agent (CA) CanAnswer to answer questions about colorectal cancer and suggest follow-up user queries. Our between-subjects study showed that compared to the condition of using online community webpages and a CA without community data, CanAnswer significantly enhanced the number of knowledge points recalled by users as well as their engagement in the learning task. This finding demonstrated the value of computationally structured community data for learning purposes (Peng et al., 2024; Wu et al., 2024) and indicated the benefits of CA for engaging users in learning tasks (Peng et al., 2024; Wambsganss et al., 2021; Ruan et al., 2019).
@@
Our computational approach and the design in CanAnswer can serve as starting points for supporting other fields of community-based knowledge acquisition, such as visual design (Peng et al., 2024), pregnancy (Wang et al., 2021), diabetes (Huh and Ackerman, 2012), anxiety (Wu et al., 2024), and so on. For example, our computational workflow, as implemented in CanAnswer, can be readily adapted to process data from online design communities for learning purposes. By extracting professional Q&A pairs (e.g., content about design principles or usability testing) from online design communities, researchers can compile a knowledge base to help the conversational agent generate reliable responses. Additionally, questions from authentic expert-novice conversations can be clustered and categorized to create suggested follow-up questions for users, assisting them in formulating better inquiries. Design-related posts from peers can be processed using a similar approach as outlined in subsection 5.2.6. Researchers can apply filtering and classification based on specific criteria (e.g., mentioning certain visual elements or UI components) to ensure the delivery of high-quality, contextually appropriate real-world cases to users. While CanAnswer is tailored for health knowledge in online communities, the features it introduces—such as suggested follow-up questions, input suggestions, and real-world case references—are broadly applicable to supporting user interaction with other types of community content. For instance, apart from creating a new post to ask for suggestions on language learning in the immigrant communities (Liang, 2024), users can ask their questions to an agent like CanAnswer, which can answer the questions based on the processed community data and guide users to learn related topics like job hunting and travel.
@@ 8.1.1. Convert Posts in Peer-led Communities to Short and Interactive Pieces
In our implementation (see Figure 2), we processed posts from RED using tag-based clustering, categorizing them into nine groups. During the Q&A process, when users explicitly requested specific examples, CanAnswer matched the context of the query to the most relevant category and retrieved the corresponding posts, presenting them in their original form. However, as illustrated in Figure 1E, some examples are lengthy and contain redundant information, which may contribute to the instances where users in our experiment perceived the provided examples as irrelevant. Similar issues appear in previous research (Peng et al., 2024; Wu et al., 2024; Peng et al., 2020) in which participants complained the lengthy posts in the community. To enhance relevance, we suggest that future work should further decompose the categorized posts e.g., based on the contained content about the symptoms, diagnosis, and recovery of a certain disease. The tool could further present such structured content within a post in an interactive manner to improve its interactivity, which is a shortcoming as indicated by two participants (Table 5). For example, it could first describe the background and the symptoms of the poster and prompt a multiple-choice question that asks the user to choose which disease the poster may have. As demonstrated by Peng et al. (2024), the answer of this question should be sourced from the original post, and the distracting options could be retrieved from other posts within the same category.
@@ 8.1.2. Enhance Peer-Contributed Community Data with Professional Content
As shown in Figure 2, we tailored our approach to the distinct characteristics of different data source by separately processing and utilizing content from doctor-led and peer-led online communities. Specifically, the disease lookup documents from Dingxiang Doctor were employed to enhance the credibility of the CanAnswer’s answers, while patient-doctor conversations were utilized to anticipate potential follow-up questions from users. Differently, posts from RED were utilized to provide concrete cases for users’ reference. Our experiment demonstrated that incorporating data from Dingxiang Doctor effectively enhanced the credibility of the agent’s responses, ensuring they were well-structured and consistent, and also facilitated a more convenient questioning process. However, both users and experts expressed concerns regarding the credibility of content sourced from RED. To address this issue, we recommend that future work should integrate peer-contributed and doctor-contributed content for Q&A. Specifically, professional documentation can be utilized to annotate and elaborate on the specialized knowledge present in peer-contributed posts, thereby guiding users to gain deeper insights from these cases (Greenhow and Galvin, 2020; McLean, 2016). This approach would leverages the complementary strengths of both sources to enhance the educational value and credibility of the information presented.
@@ 8.2. Ethical Considerations
To address this concern, future iterations of CanAnswer could integrate persuasive strategies aimed at reinforcing the importance of consulting healthcare professionals after engaging with health-related content. Experts suggested that posts detailing patients’ experiences with medical care, particularly those describing the benefits of timely diagnosis and treatment, could serve as effective prompts for encouraging users to take appropriate medical action. Beyond leveraging such community-driven narratives, future systems could also incorporate insights from research on persuasive techniques (Haque et al., 2024; Li et al., 2024b; Park et al., 2024; Ding et al., 2022). For instance, CanAnswer could simulate interactive role-playing scenarios in which users are tasked with persuading someone else to undergo a medical examination, such as a colonoscopy. This approach, inspired by prior work on persuasive role-play in health education (Hedderich et al., 2024), could reinforce users’ own motivation to pursue medical screening. Regardless of the specific strategies employed, it remains essential for such systems to transparently communicate that the presented information is for reference only and to actively guide users toward consulting medical professionals for informed decision-making.

### v2

@@ 8. Discussion
In this paper, we have designed, developed, and evaluated CanAnswer that leverages community data for supporting users in learning about colorectal designs. Our formative study reveals that online communities are important sources for people to gain informational and emotional support about certain diseases, inline with previous work that highlights the benefits of online communities (ask_the_doctor; burgess2023healthcare; emotional_support; sinha2018use; kendal2017moderated). However, the unstructured and/or unauthoritative nature of the content in these communities could be troublesome (see Section 3.4.2). We developed a computational workflow (Figure 2) to identify and structure the useful content from a professional online community Dingxiang Doctor and a peer-led community RED, which enables our conversational agent (CA) CanAnswer to answer questions about colorectal cancer and suggest follow-up user queries. Our between-subjects study showed that compared to the condition of using online community webpages and a CA without community data, CanAnswer significantly enhanced the number of knowledge points recalled by users as well as their engagement in the learning task. This finding demonstrated the value of computationally structured community data for learning purposes (DesignQuizzer; wu2024comviewer) and indicated the benefits of CA for engaging users in learning tasks (DesignQuizzer; wambsganss2021arguetutor; ruan2019quizbot).
@@
Our computational approach and the design in CanAnswer can serve as starting points for supporting other fields of community-based knowledge acquisition, such as visual design (DesignQuizzer), pregnancy (wang2021cass), diabetes (online_diabete_community), anxiety (wu2024comviewer), and so on. For example, our computational workflow, as implemented in CanAnswer, can be readily adapted to process data from online design communities for learning purposes. By extracting professional Q&A pairs (e.g., content about design principles or usability testing) from online design communities, researchers can compile a knowledge base to help the conversational agent generate reliable responses. Additionally, questions from authentic expert-novice conversations can be clustered and categorized to create suggested follow-up questions for users, assisting them in formulating better inquiries. Design-related posts from peers can be processed using a similar approach as outlined in subsection 5.2.6. Researchers can apply filtering and classification based on specific criteria (e.g., mentioning certain visual elements or UI components) to ensure the delivery of high-quality, contextually appropriate real-world cases to users. While CanAnswer is tailored for health knowledge in online communities, the features it introduces—such as suggested follow-up questions, input suggestions, and real-world case references—are broadly applicable to supporting user interaction with other types of community content. For instance, apart from creating a new post to ask for suggestions on language learning in the immigrant communities (liang2024immigration), users can ask their questions to an agent like CanAnswer, which can answer the questions based on the processed community data and guide users to learn related topics like job hunting and travel.
@@ 8.1.1. Convert Posts in Peer-led Communities to Short and Interactive Pieces
In our implementation (see Figure 2), we processed posts from RED using tag-based clustering, categorizing them into nine groups. During the Q&A process, when users explicitly requested specific examples, CanAnswer matched the context of the query to the most relevant category and retrieved the corresponding posts, presenting them in their original form. However, as illustrated in Figure 1E, some examples are lengthy and contain redundant information, which may contribute to the instances where users in our experiment perceived the provided examples as irrelevant. Similar issues appear in previous research (DesignQuizzer; wu2024comviewer; emotional_informational2) in which participants complained the lengthy posts in the community. To enhance relevance, we suggest that future work should further decompose the categorized posts e.g., based on the contained content about the symptoms, diagnosis, and recovery of a certain disease. The tool could further present such structured content within a post in an interactive manner to improve its interactivity, which is a shortcoming as indicated by two participants (Table 5). For example, it could first describe the background and the symptoms of the poster and prompt a multiple-choice question that asks the user to choose which disease the poster may have. As demonstrated by DesignQuizzer, the answer of this question should be sourced from the original post, and the distracting options could be retrieved from other posts within the same category.
@@ 8.1.2. Enhance Peer-Contributed Community Data with Professional Content
As shown in Figure 2, we tailored our approach to the distinct characteristics of different data source by separately processing and utilizing content from doctor-led and peer-led online communities. Specifically, the disease lookup documents from Dingxiang Doctor were employed to enhance the credibility of the CanAnswer’s answers, while patient-doctor conversations were utilized to anticipate potential follow-up questions from users. Differently, posts from RED were utilized to provide concrete cases for users’ reference. Our experiment demonstrated that incorporating data from Dingxiang Doctor effectively enhanced the credibility of the agent’s responses, ensuring they were well-structured and consistent, and also facilitated a more convenient questioning process. However, both users and experts expressed concerns regarding the credibility of content sourced from RED. To address this issue, we recommend that future work should integrate peer-contributed and doctor-contributed content for Q&A. Specifically, professional documentation can be utilized to annotate and elaborate on the specialized knowledge present in peer-contributed posts, thereby guiding users to gain deeper insights from these cases (greenhow2020teaching; mclean2016case). This approach would leverages the complementary strengths of both sources to enhance the educational value and credibility of the information presented.
@@ 8.2. Ethical Considerations
To address this concern, future iterations of CanAnswer could integrate persuasive strategies aimed at reinforcing the importance of consulting healthcare professionals after engaging with health-related content. Experts suggested that posts detailing patients’ experiences with medical care, particularly those describing the benefits of timely diagnosis and treatment, could serve as effective prompts for encouraging users to take appropriate medical action. Beyond leveraging such community-driven narratives, future systems could also incorporate insights from research on persuasive techniques (haque2024butt; StayFocused; ADHD_child; ding2022talktive). For instance, CanAnswer could simulate interactive role-playing scenarios in which users are tasked with persuading someone else to undergo a medical examination, such as a colonoscopy. This approach, inspired by prior work on persuasive role-play in health education (bully_roleplay_education), could reinforce users’ own motivation to pursue medical screening. Regardless of the specific strategies employed, it remains essential for such systems to transparently communicate that the presented information is for reference only and to actively guide users toward consulting medical professionals for informed decision-making.